College of the Holy Cross, Fall Semester, 2007

MATH 131, Section 01, Final Exam

Saturday, December 8, 2:30 PM Professor Little

Your Name:

Directions. You must show all work required to derive each answer for full credit. Please place final answers in the boxes provided. Use the back of the preceding page if you need more space for scratch work. There are 200 possible points, distributed as shown below.

Please do not write in the space below

Problem Points/Poss
1  / 20
2  / 25
3  / 15
4  / 25
5  / 15
6  / 25
7  / 25
8  / 20
9  / 20
10  / 10
Total  /200

ℋ 𝒜𝒫𝒫𝒴 ℋ 𝒪ℒℐ𝒟𝒜𝒴𝒮 !

MATH 131-01, Final Exam 2 1. [5 points each] Circle the number of the graph showing each of the following functions. Note that there is one more graph than there are formulas.

(a) f(x) = 2 − 2 sin(πx) I II III IV V (b) f(x) = \frac{x^{2}}{x^{2} − 1} I II III IV V (c) f(x) = ln(\sqrt{x}) I II III IV V (d) f(x) = e^{x} + e^{−x} I II III IV V

I: II: III: IV: